# PhreshPhish: A Real-World, High-Quality, Large-Scale Phishing Website Dataset and Benchmark

### v1

@@ E.1 Setup, Hyperparameters and Training Details
Linear model We implemented a linear support vector classifier (using LinearSVC) with squared hinge loss and L2 regularization. Key hyperparameters included: tolerance for convergence set to $310^{-}$ 3, dual set to true – due to large number of features, and class_weight set to ’balanced’ – to weight samples inversely proportional to their class frequencies. The model used a combination of a few tens of hand-crafted features and a large number of n-gram features. The hand-crafted features included URL-based, such as the overall length of the URL, characteristics of the path, domain and subdomain structure, and lengths of individual components; and those derived from the HTML DOM structure, such as the number of nodes and edges in the DOM tree, to capture the structural complexity of the web page. The training converged in $\sim$ 24 epochs.
@@
FNN model We used two types of n-gram features as input to our feedforward neural network (FNN) model. The first type consists of character-level 2-, 3-, and 4-grams extracted from the URL and selected sections of the parsed HTML, including: visible text, HTML tags, scripts, links, image sources, and embedded image data. The second type of n-gram feature is derived from XPath sequences extracted from the HTML DOM. Unlike the character-based n-grams, these XPath n-grams are computed at the tag level. For example, the 2-gram representation of the XPath sequence <document>/<div>/<p>/<a> includes the segments: <document>/<div>, <div>/<p>, and <p>/<a>. Each character-level or XPath n-gram is treated as an independent binary feature for input to the FNN.
The feature extraction process initially produced approximately 15.3 million binary features. We selected the 300,000 most frequent features from the training set for use in model training. Training was conducted over 10 epochs. The entire pipeline, including feature extraction and model training, completed in approximately 32 hours.
@@
BERT-based model We fine-tuned gte-large pre-trained embedding model on our training dataset using Hugging Face trainer with adamw optimizer, a learning rate of $10^{-5}$, momentum (adam_beta1) of 0.9 and a weight decay of $10^{-4}$. The model was fine-tuned for 2 epochs with auto_find_batch_size set to true for auto batch size discovery, which resulted in a batch size of 8 per device (32 in all). For the 2 epochs, it took about 27 hours.
@@
LLM-based prediction We used OpenAI’s gpt-4o-mini model in batch mode to classify webpages as phishing or benign based on the input URL and HTML content. A system prompt was crafted to instruct the model to assess social engineering indicators and output a JSON containing two fields: a phishing flag (boolean) and a phishing risk score (1–10). The HTML input is preprocessed to remove non-essential elements and truncated to stay within token limits. Each request includes the URL and simplified HTML in the prompt, and the model returns structured output used for phishing detection analysis. The token limit is set to 2500 for HTML and 500 for the prompt template. No token limit was imposed on the URL.
HTML simplification
@@
To help shorten the input text, we remove unnecessary HTML elements such as metadata and styling. Following guidance from prior work in phishing detection using LLMs [26], the simplification process consists of: (1) Removing comments, <style>, and <script> tags; (2) Preserving content tags: p, a, img, h1--h6, ul, ol, li; (3) Filtering out tags with no text content; (4) Normalizing links and images by stripping http(s):// and www.; (5) If token limit is exceeded despite above steps, HTML elements are trimmed from the center until token limit is met.
System prompt
@@
"You are an expert in analyzing URLs and multilingual HTML to classify webpages as phishing or legitimate. Focus only on identifying whether the page is phishing and its corresponding risk score. First perform thorough reasoning on the URL and multi-lingual HTML and then classify with a confidence score. Given the following URL and HTML, analyze for social engineering techniques commonly used in phishing attacks: 1. PHISHING: Is the webpage phishing? (True if phishing, False otherwise) 2. SCORE: Phishing risk on a 1-10 scale (1 = least likely, 10 = most likely). Return the result in JSON format with the following fields: ‘phishing‘: boolean, ‘score‘: integer"
E.2 Evaluation results
Table 8 shows the mean average precision and the mean precision at recall$=0.9$ of the four models on the test set (base rate: 18.3%) [^6] This is same as the last column of Table 2 but with confidence intervals added. The means here may not exactly match the ones in the other table since these were computed from bootstrap samples. The 95% confidence intervals are computed based on 1000 stratified bootstrap samples on the test set.
Test set (base rate: 18.3%)
Model Mean Lower Upper
Linear AP 0.9815 0.9795 0.9835
P@R = 0.90 0.9672 0.9596 0.9734
FNN AP 0.9339 0.9296 0.9383
P@R = 0.90 0.7991 0.7805 0.8141
GTE AP 0.9918 0.9906 0.9929
P@R = 0.90 0.9855 0.9825 0.9888
LLM AP 0.8368 0.8280 0.8453
P@R = 0.90 0.6930 0.6217 0.7771
Table 8: Mean average precision (AP) and precision at 90% recall for the test dataset, with the corresponding 95% confidence interval bounds provided in the ’Lower’ and ’Upper’ columns.
@@ E.2 Evaluation results
Next, we provide several visualizations to show the performance of the four models on the test and the benchmark datasets.
@@
Prediction Score Histograms Figure 10 shows the score histogram for each of the models, with separate plots for benign and phishing points. The GTE model which has the best performance based on average precision also shows good class separation, with minimal overlap in the central score region. Further, the sharp peaks near zero and one, for the benign and phishing points, respectively, shows high confidence in correctly classified instances. Other models show multimodal distributions for the two classes, with lesser separation and greater overlap in the mid-score region.
@@
ROC and PR curves Figure 11 presents the ROC and precision-recall (PR) curves for the test dataset. Most models show strong performance, with the exception of the LLM-based model. The ROC-AUC values are quite close to 1.0, since the FPR values are numerically low, even at 18% base rate. As the base rate decreases, the FPP for a given level of precision tends to become even smaller. Consequently, we omit ROC curves for lower base rates.
@@
Confusion matrices Figure 12 shows the trade off between precision and recall as the model threshold is changed. As the threshold increases the recall decreases while the precision increases. For example, at threhold of 0.7, the GTE model achieves a precision of 98.9% at a recall of 88%.
PR curves on the Benchmark datasets The precision-recall (PR) curves in Figure 13 show how the performance of the models decreases with decrease in base rate. Each plot presents five PR curves for each of the five base rates (0.05%, 0.1%, 0.5%, 1% and 5%).
@@
Figure 10: Model score distribution on the test dataset.
@@
Figure 11: ROC and PR curves for the test dataset.
@@
Figure 12: Confusion matrices at different thresholds (t) for the test dataset. In addition, the precision (P) and recall (R) at that threshold is also reported.
@@
Figure 13: Precision-recall (PR) curves for the benchmark datasets are shown at five distinct base rates. For each base rate, the dark curve represents the average PR curve.

### v2

@@ C.1. Setup, Hyperparameters and Training Details
Linear model We implemented a linear support vector classifier (using LinearSVC in scikit-learn) with squared hinge loss and L2 regularization. Key hyperparameters included: tolerance for convergence set to $10^{-3}$, dual=True and class_weight=’balanced’. The model used a combination of a few tens of hand-crafted features and a large number of n-gram features. The hand-crafted features included URL-based, such as the overall length of the URL, characteristics of the path, domain and subdomain structure, and lengths of individual components; and those derived from the HTML DOM structure, such as the number of nodes and edges in the DOM tree, to capture the structural complexity of the web page. The training converged in 24 epochs.
@@
FFN model We used two types of n-gram features as input to our feedforward neural network (FFN) model. The first type consists of character-level 2-, 3-, and 4-grams extracted from the URL and selected sections of the parsed HTML, including: visible text, HTML tags, scripts, links, image sources, and embedded image data. The second type of n-gram feature is derived from XPath sequences extracted from the HTML DOM. Unlike the character-based n-grams, these XPath n-grams are computed at the tag level. For example, the 2-gram representation of the XPath sequence <document>/<div>/<p>/<a> includes the segments: <document>/<div>, <div>/<p>, and <p>/<a>. Each character-level or XPath n-gram is treated as an independent binary feature for input to the FFN. The feature extraction process initially produced approximately 66.5 million binary features. We selected the 1.6 million most frequent features from the training set for use in model training. Training was conducted over 20 epochs. The entire pipeline, including feature extraction and model training, completed in approximately 45 hours.
@@
BERT-based model We fine-tuned gte-large pre-trained embedding model on our training dataset using Hugging Face trainer with adamw optimizer, a learning rate of $10^{-5}$, momentum of 0.9 and a weight decay of $10^{-4}$. The model was fine-tuned for 2 epochs, taking approximately about 27 hours.
@@
LLM-based prediction We used Google’s gemini-2.5-pro model in batch mode to classify webpages as phishing or benign based on the input URL and HTML content. A system prompt was crafted to instruct the model to assess social engineering indicators and output a JSON containing two fields: a phishing flag (boolean) and a phishing risk score (1–10). The HTML input is preprocessed to remove non-essential elements and truncated to stay within token limits. Each request includes the URL and simplified HTML in the prompt, and the model returns structured output used for phishing detection analysis.
@@
HTML simplification: To help shorten the input text, we remove unnecessary HTML elements such as metadata and styling. Following guidance from prior work in phishing detection using LLMs [koide2024chatphishdetector], the simplification process consists of: (1) Removing comments, <style>, and <script> tags; (2) Preserving content tags: p, a, img, h1--h6, ul, ol, li; (3) Filtering out tags with no text content; (4) Normalizing links and images by stripping http(s):// and www.; (5) If token limit is exceeded despite above steps, HTML elements are trimmed from the center until token limit is met.
System prompt:
@@
      You are an expert in analyzing URL and multi-lingual HTML    to classify webpage as phishing or legitimate page.    Focus only on identifying if the page is phishing and    its corresponding risk score. Given the URL and HTML,    perform the following analysis for any Social Engineering    techniques often used in phishing attacks:      First perform thorough REASONING on the URL and HTML    looking for suspicious elements or inconsistencies,    domain issues, detect content red flags, evaluate    security indicators and inconsistencies, and then    classify with a confidence score.      1. **PHISHING:** Whether the webpage is a phishing page    or a legitimate page. (True if phishing, False if    legitimate)    2. **SCORE:** Indicates phishing risk on a scale of    0.0 to 1.0 (inclusive), where 0.0 is the least likely    to have phishing content and 1.0 is the most likely.    3. **REASON:** Briefly (one sentence) explain the    reasoning behind your determination.      Provide the extracted information in JSON format with the    following keys:    - phishing: boolean (whether the site is a phishing site    or a legitimate site)    - score: float (indicates phishing risk on a scale of 0.0    to 1.0)    - reason: str (one sentence reason)
We note that Gemini refused to score 52 samples (46 benign and 6 phishing) due to containing prohibited contents. These instances were exceluded from the analysis when computing the performance of this model.
C.2. Benchmark evaluation results
@@
We provide several visualizations to show the performance of the four models on the full (24%) benchmark dataset. Similar plots are also shown for the test set in C.3
@@
Prediction score histograms Figure 11 (left) shows the score histogram on the benchmark set for each of the models, with separate plots for benign and phishing points. The GTE model which has the best performance based on average precision also shows good class separation, with minimal overlap in the central score region. Further, the sharp peaks near zero and one, for the benign and phishing points, respectively, shows high confidence in correctly classified instances. Other models show multimodal distributions for the two classes, with lesser separation and greater overlap in the mid-score region.
@@
ROC and PR curves Figure 11 (right) presents the ROC and precision-recall (PR) curves for the benchmark dataset. Most models show strong performance, with the exception of the LLM-based model. The ROC-AUC values are quite close to 1.0, since the FPR values are numerically low, even at 24% base rate. As the base rate decreases, the FPR for a given level of precision tends to become even smaller. Consequently, we omit ROC curves for lower base rates.
@@
Confusion matrices Figure 12 shows the trade off between precision and recall as the model threshold is changed. As the threshold increases the recall decreases while the precision increases. For example, at threhold of 0.7, the GTE model achieves a precision of 99.6% at a recall of 80.3%.
@@
Figure 11. Model performance on the full (24%) benchmark dataset: (left) score distributions, (right) ROC and PR curves.
@@
Figure 12. Confusion matrices at different thresholds (t) for the full (24%) benchmark dataset. In addition, the precision (P) and recall (R) at that threshold is also reported.
C.3. Test evaluation results
@@
Figure 13. Model performance on the test dataset: (left) score distributions, (right) ROC and PR curves.
@@
Figure 14. Confusion matrices at different thresholds (t) for the test dataset. In addition, the precision (P) and recall (R) at that threshold is also reported.